OpenStudy (anonymous):

H.E.L.P M.E ! WHO EVER HELPS GETS MEDAL AND FAN ! The Number of red frogs exceeded the number of blue frogs by 80. The number of green frogs was 20 less than the number of blue frogs. If there were 120 blue frogs, what was the sum of the reds, the blues and the greens? IM SORRY FOR THIS LONG QUESTION !

OpenStudy (igreen):

We can make equations with the information given. The Number of red frogs exceeded the number of blue frogs by 80. \(\sf r + 80 = b\) The number of green frogs was 20 less than the number of blue frogs. \(\sf g - 20 = b\) there were 120 blue frogs \(\sf b = 120\) Plug in 120 into the first two equations for \(\sf b\) and solve for \(\sf r\) and \(\sf g\).

OpenStudy (anonymous):

Oh, So If I find what ( R , G , AND B) Are , That would be the answer to this question ?

OpenStudy (anonymous):

i mean r and g

OpenStudy (igreen):

\(\sf b\) is already given..and we have two equations to find \(\sf r \) and \( \sf g\), so plug in \(\sf 120\) for \(\sf b\) into those equations to find \(\sf r \) and \( \sf g\). Then after you find all 3 of them, add them up.

OpenStudy (igreen):

Oops, I messed these equations up.. \(\sf r = b + 80\) \(\sf r = 120 + 80\) \(\sf g = b - 20\) \(\sf g = 120 - 20\)

OpenStudy (anonymous):

OKAY THANKS SO MUCH FOR YOUR HELP !!!!!

OpenStudy (igreen):

No problem.